OpenStudy (blu.acuna):

rewrite 32,000,000 in the power of ten

OpenStudy (sooobored):

in scientific notation? 32,000,000 is equivalent to 32* 100,000 or 3.2*1,000,000

OpenStudy (sooobored):

and 100,000 is equal to 10^5 and 1,000,000 is equal to 10^6 scientific notation would be the 3.2 *10^...

OpenStudy (p_sid):

32*10^6=3.2*10^7=0.32*10^8